OpenStudy (cedric):

Your piggy bank has a total of 47 coins in it; some are dimes and some are nickels. If you have a total of $3.95, how many nickels do you have?

OpenStudy (cedric):

HELP! ! ! !

OpenStudy (ajprincess):

Let the number of dimes be x and number of nickels be y. x+y=47-(1) 10x+5y=395-(2) (1)*10=>10x+10y=470-(3) (3)-(2)=>5y=75 Solve for y.

OpenStudy (cedric):

15?